Will NicOs precipitate from a solution that initially contains 0.0020 M NazCOs? 0.150 M NiCl2 and 1.20 M NH3? 9.
Answer #1

Ans. Ksp of NiCO3 = 1.42 * 10^ -7

Na2CO3 = 2Na+ + CO3 2-

2 * 0.002 0.002

NiCl2 = Ni 2+ + 2Cl-

0.15 2 * 0.15

[CO32-] = 0.002 M

[Ni2+] = 0.15

[Ni2+] [CO3 2-] = 0.15 * 0.002 = 3.0 * 10^-4 M^2

ionic product of NiCO3 is more than that of Ksp and hence it will precipitated